Fore Farms reported a pretax operating loss of $268 million for financial reporting purposes in 2021. Contributing to the loss were (a) a penalty of $4 million assessed by the Environmental Protection Agency for violation of a federal law and paid in 2021 and (b) an estimated loss of $20 million from accruing a loss contingency. The loss will be tax deductible when paid in 2022.

The enacted tax rate is 25%. There were no temporary differences at the beginning of the year and none originating in 2021 other than those described above. Taxable income in Fores’s two previous years of operation was as follows:

2019 $ 120 million
2020 $ 96 million


Required:
1. Prepare the journal entry to recognize the income tax benefit of the net operating loss in 2021. Assume Fore will carry back its NOL to prior years.
2. What is the net operating loss reported in 2021 income statement?
3. Prepare the journal entry to record income taxes in 2022 assuming pretax accounting income is $288 million. No additional temporary differences originate in 2022.
Answer #1

1 Particulars DR CR Receivable-Income tax refund $ 54.00 =(120+96)*25% Deferred tax asset $ 12.00 =(268-120-96-4) *25% Income